maya, rafael, and carlos have a total of $90 in their wallet. Maya has $5 more than carlos. rafael has 3x what carlos has. how much does each has
Answer by rfer(16322)

Carlos=x
Maya=x+5
Rafael=3x
---------------
x+x+5+3x=90
5x=90-5
x=85/5
x=17
x+5=22
3x=51